OpenStudy (falling_in_katt):

Simplify the expression where possible. \[(-4x ^{2})^{2}\]

OpenStudy (abdullahm):

Hint: \(\sf\Large (a^b)^c = a^{b \times c}\)

OpenStudy (jdoe0001):

ditto

OpenStudy (falling_in_katt):

\[-4^{4}\]?

OpenStudy (danjs):

\[[-4^1*x^2]^2\]

OpenStudy (falling_in_katt):

What? I'm confused.

OpenStudy (danjs):

\[= (-4^1)^2*(x^2)^2\]

OpenStudy (abdullahm):

Another hint: \(\sf\Large (xy)^z = x^{z} y^{z}\)

OpenStudy (danjs):

@AbdullahM wrote the correct general power rules to remember... good luck

OpenStudy (abdullahm):

(:

OpenStudy (falling_in_katt):

\[-4x ^{4}?\]

OpenStudy (jdoe0001):

\(\bf (-4x ^{2})^{2}\implies (-4^1x^{2})^2\implies (-4^1)^2(x^2)^2\implies -4^{1\cdot 2}x^{2\cdot 2}\implies -4^2x^4\)

jimthompson5910 (jim_thompson5910):

@Falling_In_Katt when you square the -4, it turns into (-4)^2 = (-4)*(-4) = +16 notice how it's positive and not negative

OpenStudy (falling_in_katt):

Ohhh Thank you!

jimthompson5910 (jim_thompson5910):

It's positive because two negatives multiply to a positive

OpenStudy (jdoe0001):

hmmm actually shold be positive yes

OpenStudy (jdoe0001):

\(\bf (-4x ^{2})^{2}\implies (-4^1x^{2})^2\implies (-4^1)^2(x^2)^2 \\ \quad \\ (-4)(-4)(x)^4\implies +16x^4\)